One number is 1/5 of another number. The sum of the two numbers is 12. Find the two numbers. Use a comma to separate your answers.
Found 2 solutions by ikleyn, amfagge92:
Answer by ikleyn(52797)
Answer by amfagge92(93)

let x and y be the numbers
x=1/5*y...eqn i
x+y=12....eqn ii
subtitute x in eqn ii
y/5+y=12
6y=60
y=10
subtitute y in eqn....i
x=10/5
=2
therefore 2,10 are the numbers.
N.B the first answer from Madam ikleyn is absolutely correct! i just elucidate.